What's (a)2v + u, (b)u -3v, and (c)5u - v in the equation of u = i - j, v = 6i + 9j?
Have, u=(1,-1), v=(6,9)
2v + u =2(6,9) + (1,-1) = (12,18) + (1,-1) = (12+1,18-1) = (13,17)
u - 3v =(1,-1) -3(6,9) = (1,-1) - (18,27) = (1-18,-1-27) = (-17,-28)
5u - v = 5(1,-1) - (6,9) = (5,-5) - (6,9) = (5-6,-5-9) = (-1,-14)
but need to label (a), (b), and (c).

Answer by Alan3354(69443)

What's (a)2v + u, (b)u -3v, and (c)5u - v in the equation of u = i - j, v = 6i + 9j?
Have, u=(1,-1), v=(6,9)
2v + u =2(6,9) + (1,-1) = (12,18) + (1,-1) = (12+1,18-1) = (13,17)
u - 3v =(1,-1) -3(6,9) = (1,-1) - (18,27) = (1-18,-1-27) = (-17,-28)
5u - v = 5(1,-1) - (6,9) = (5,-5) - (6,9) = (5-6,-5-9) = (-1,-14)
but need to label (a), (b), and (c).
--------------
(a) 13i + 17j
(b) -17i - 28j
(c) -i - 14j